### v1

# Self-Modeling Robots by Photographing

Abstract
Self-modeling enables robots to build task-agnostic models of their morphology and kinematics based on data that can be automatically collected, with minimal human intervention and prior information, thereby enhancing machine intelligence. Recent research has highlighted the potential of data-driven technology in modeling the morphology and kinematics of robots. However, existing self-modeling methods suffer from either low modeling quality or excessive data acquisition costs. Beyond morphology and kinematics, texture is also a crucial component of robots, which is challenging to model and remains unexplored. In this work, a high-quality, texture-aware, and link-level method is proposed for robot self-modeling. We utilize three-dimensional (3D) Gaussians to represent the static morphology and texture of robots, and cluster the 3D Gaussians to construct neural ellipsoid bones, whose deformations are controlled by the transformation matrices generated by a kinematic neural network. The 3D Gaussians and kinematic neural network are trained using data pairs composed of joint angles, camera parameters and multi-view images without depth information. By feeding the kinematic neural network with joint angles, we can utilize the well-trained model to describe the corresponding morphology, kinematics and texture of robots at the link level, and render robot images from different perspectives with the aid of 3D Gaussian splatting. Furthermore, we demonstrate that the established model can be exploited to perform downstream tasks such as motion planning and inverse kinematics.
keywords:
Self-Model, 3D Gaussian Splatting, Dynamic Reconstruction
1 Introduction
High-quality models are essential for effective robot planning, control, and visualization. Traditional modeling approaches demand substantial human effort to gather the necessary parameter information of the robot, and over time, the quality of these models tends to deteriorate due to various factors, such as unexpected damage, joint malfunctions, and wear. Therefore, the development of methods that enable robots to collect data and learn their own models autonomously is critical for ensuring the successful execution of tasks and advancing the level of machine intelligence. Inspired by human behavior (Rochat, 2003), the concept of robot self-modeling has recently gained attention among researchers (Chen et al., 2022; Díaz Ledezma and Haddadin, 2023). A robot self-model is a general-purpose representation of a robot’s physical shape and structure that can be acquired at any time without human intervention (Kwiatkowski, 2022). Through self-modeling, robots can autonomously and continuously update their knowledge of both morphology and kinematics, which not only ensures that the model information remains up to date and adaptable to various unforeseen circumstances but also simplifies the modeling process for newly designed robots.
Early approaches focused on designing task-specific self-modeling techniques, (Kwiatkowski and Lipson, 2019; Chen et al., 2021; Hang et al., 2021) such as modeling end-effector positions (Mathew et al., 2014), tilt angles (Bongard et al., 2006), or joint velocities (Sanchez-Gonzalez et al., 2018) according to task requirements. However, due to the lack of generality and adaptability in these methods, recent work has shifted toward modeling the whole robot to construct task-agnostic self-models. For instance, Chen et al. (2022) employed visual data with depth information to enable self-modeling of the overall morphology of a 4-degree-of-freedom robot, which was later validated on robots with higher degrees of freedom by Yang et al. (2024a). Their approach allows the robot to determine its spatial occupancy based on input joint configurations but cannot directly compute the kinematic information (e.g., position) of specific links. To address link-level modeling,  Yang et al. (2024a) developed multiple independent models for each robot link and integrated them into a comprehensive self-model. However, these methods require depth information to create self-models, which necessitates expensive devices (e.g., depth cameras and LiDAR) and is hard to obtain in practice. In addition to visual data, inertial measurement unit (IMU) data can also be utilized for the self-modeling of robot (Díaz Ledezma and Haddadin, 2023), which necessitates the attachment of numerous IMU sensors to the robot links. Notably, despite the considerable advancements in robot self-modeling concerning morphology and kinematics, the aforementioned approaches have neglected the modeling of texture features (such as surface color and patterns), which are crucial for the visualization of robots in simulation environments and digital twin systems. Thus, beyond morphology and kinematics, texture modeling remains a significant challenge in achieving comprehensive robot self-modeling.
Given the challenges and high costs of data acquisition in the previously discussed self-modeling methods and the importance of texture modeling, a visual self-modeling approach utilizing only two-dimensional (2D) images presents a more efficient and cost-effective alternative. In this paper, we develop a scheme based on 3DGS (Kerbl et al., 2023) and neural networks for modeling the robot’s morphology, kinematics, and texture. 3DGS is a newly developed technology for 3D scene reconstruction, which represents 3D scenes with a large number of optimizable, flexible, and expressive 3D Gaussian ellipsoids, enabling real-time rendering of scene images. With the assistance of 3DGS, we can acquire 3D information, which is necessary for robot self-modeling from 2D images. An overview of our method is demonstrated in Figure1a. To optimize the 3D Gaussians and kinematic network, the robot only needs to gather multi-view RGB images under various configurations, along with corresponding joint configuration data and camera parameters. The static morphology and texture information of the robot’s links are embedded in the 3D Gaussians, each of which is an ellipsoid defined by its position, covariance matrix, color, and opacity. By clustering these 3D Gaussians, the bones of the robot model are constructed, with transformations of the bones determined by the kinematic network. By feeding the kinematic neural network with joint angles, we can utilize the well-trained self-model to describe the corresponding morphology, kinematics, and texture of robots at the link level and render robot images from different perspectives with the aid of 3DGS. Besides, the self-model can be employed to directly control the robot to perform downstream tasks such as motion planning and inverse kinematics. As illustrated in Figure1b, compared to previous methods, the proposed method relies solely on easily accessible data (e.g., multi-view RGB images) for learning while producing a high-quality, texture-aware, and link-level self-model of robots.
Figure 1: Robot self-modeling from images. a Given a robot whose model is unknown, our method can build a self-model of the robot from multiple RGB images. We capture multi-view photos of the robot at different joint configurations, and train the 3D Gaussians and the kinematic network, which are responsible for representing the static morphology and texture of robots and neural bone deformation, respectively. The well-trained model can be utilized to render robot images with the aid of 3D Gaussian splatting and perform downstream tasks. b Compared to the methods proposed by Chen et al. (2022) and Schulze and Lipson (2024), our method requires no depth information and can establish a high-quality, texture-aware and link-level self-model.
Overall, the contributions of this paper are:
  • We propose a robot self-modeling method that can be built from images without 3D information based on 3DGS. The robot self-model can correctly reconstruct the robot’s morphology, kinematics, and texture.
  • We model the robot at the link level. The self-model automatically identifies different links of the robot and can predict their movement separately.
  • We use the self-model to control the robot. Experiments show that the self-model can control the robot to accomplish downstream tasks without additional training.
2 Related work
The main challenge of self-modeling robots from 2D visual input is how to reconstruct 3D models from 2D images or videos. Currently, there are two solutions: Neural Radiance Field (Mildenhall et al., 2021) and 3D Gaussian Splatting (Kerbl et al., 2023).
2.1 3D reconstruction with NeRF
Neural Radiance Field (NeRF) has shown remarkable results in reconstructing scenes from multiple-view photos. Pumarola et al. (2021) proposed the first method to model dynamic scenes with NeRF. To accelerate the training of NeRF, methods such as (Fang et al., 2022; Müller et al., 2022) are proposed. These works motivate researchers to use dynamic NeRF to reconstruct and animate the human body (Noguchi et al., 2021)and animals (Yang et al., 2022). Noguchi et al. (2022) combined NeRF with signed-distance functions (SDFs) to learn both the appearance and the structure of articulated objects, including robots and human bodies, by observing them move from multiple views. They modeled articulated objects with several ellipsoids. They could repose articulated objects by transforming ellipsoids. Uzolas et al. (2024) extracted point clouds of articulated objects from pre-trained NeRF models and used Linear Blend Skinning (LBS) to learn Dynamic NeRF and associated skeletal models jointly. Although these works achieved modeling robots from images or videos, they did not associate robots’ joints with the built morphology models. In other words, they did not model the robots’ kinematics.
2.2 3D reconstruction with 3DGS
Though NeRF shows promising results in 3D reconstruction, it has a significant drawback. To get a high-quality scene reconstruction, it may take a long time. The 3DGS method is designed to solve this problem and can quickly complete training and build high-quality models. Similar to NeRF, researchers extended 3DGS to the reconstruction of dynamic 3D scenes (Wu et al., 2024; Yang et al., 2024c; Liang et al., 2023; Yang et al., 2024b; Lin et al., 2024).  Hu et al. (2024); Qian et al. (2024) proposed to model 3D human avatars with 3DGS. Several other approaches aim to design controllable and editable models with 3DGS. Yu et al. (2024) used pre-defined masks to choose moving parts of the modeled scene. These masks are used to control the model, and only the chosen part can be controlled. Huang et al. (2024) used control points to control the dynamic 3DGS model. Control points are automatically generated during training, which then control the movement of the whole 3DGS model with Linear Blend Skinning. Zhang et al. (2024) proposed to use neural bones to control the 3DGS model, which was similar to the method proposed by Yang et al. (2022). A neural bone is an ellipsoid parameterized by position and radius. When controlling the 3DGS model, neural bones work in the same way as real bones in human and animal bodies.
2.3 Modeling robots with 2D input
Several studies have explored using 2D images for robot self-modeling (Hu et al., 2023; Marques Monteiro et al., 2024). However, these approaches are limited to rendering images of the robot in various states and do not succeed in constructing a comprehensive model of the robot’s morphology. Schulze and Lipson (2024) was the first to solve the robot self-modeling problem using dynamic NeRF and associated the built model with input robot joints. However, NeRF-based self-modeling requires optimizing a dense multi-layer perceptron (MLP) for learning both morphology and kinematics, and uses volumetric ray-marching to render robot images. As a result, NeRF-based self-modeling suffers from problems such as difficulties in model training, high rendering costs, and suboptimal modeling quality. Lou et al. (2024) used a hybrid representation that combines 3DGS and mesh to create digital assets of robots. The model of robots can be controlled by robot joints. However, real Denavit-Hartenberg parameters are required to control the model based on joints, which requires human intervention and fails to model robot kinematics automatically.
3 Method
Figure 2: An overview of our method. We first train a static 3DGS self-model with images of the robot at zeor pose from various views. Then, we initialize neural ellipsoid bones to model the robot kinematic chain. Using images of the robot at different joint configurations, we train a dynamic self-model. The neural bones are deformed, and the relative 3D Gaussian is controlled by LBS. Finally, the model is rendered into images and compared with the ground truth images.
We propose a novel robot self-modeling method that begins by modeling the robot in a rest pose, which we define as canonical space. The robot is then transformed into various poses, which form the observation space. To implicitly represent the robot’s self-model, we utilize 3D Gaussian splatting (3DGS). The model can be deformed and rendered into images, allowing us to visualize the robot’s morphology from different perspectives. An overview of our method is provided in Figure 2.
3.1 Preliminaries
3DGS uses 3D Gaussians to represent a static scene (Kerbl et al., 2023). A 3D Gaussian is a small ellipsoid with feature parameters. Typically, a static 3DGS model consists of 10 thousands of 3D Gaussians. Each 3D Gaussian is defined by a full 3D covariance matrix $\Sigma$:
$G(x)\ =\ e^{-\frac{1}{2}{(x)}^{T}\Sigma^{-1}(x)}.$ (1)
The covariance matrix $\Sigma$ can be calculated by the rotation matrix $R$ and scaling matrix $S$ as $\Sigma=RSS^{T}R^{T}$. The rotation matrix is derived from a quaternion $q$. Each Gaussian has an opacity value $\sigma$ and a spherical harmonic coefficient $\xi$ to create view-dependent color. When rendering images, 3D Gaussians are projected to a 2D plane. Given a viewing transformation $W$, the covariance matrix $\Sigma^{\prime}$ in camera coordinates is given by:
$\Sigma^{\prime}=(JW\Sigma W^{T}J^{T})$ (2)
where $J$ is the Jacobian of the affine approximation of the projective transformation. The color $C$ of a pixel is computed by alpha blending 3D Gaussians that overlap on the pixel sorted by depth, and the color $C$ is given by:
$C=\sum_{i\in N}{c_{i}\alpha_{i}\prod_{j=1}^{i-1}{(1-\alpha_{j})}}$ (3)
where $N$ is the number of Gaussians overlapping the pixel, $c_{i}$ is the color of Gaussian $i$ calculated by spherical harmonic coefficient, $\alpha$ is given by evaluating the projected Gaussian with covariance matrix $\Sigma^{\prime}$ and multiplied with its opacity value $\sigma$. 3D Gaussian parameters $\mu,q,S,\sigma,\xi$ are optimized during training, and the number of 3D Gaussians is adaptively controlled.
3.2 Static reconstruction and neural bone
Our self-modeling method starts with modeling the robot in a rest pose. Given a set of images of a robot captured from different view angles, whose joints are all set to 0, we first reconstruct a static self-model of the robot. We then extract neural bones from the static self-model.
Inspired by BANMo (Yang et al., 2022) and BAGS (Zhang et al., 2024), we use a set of neural bones to model the kinematic structure of the robot and control the transformation of the self-model. Each bone is represented by an ellipsoid, which is defined by a center point $p\in\mathbb{R}^{3}$, a radii scaling matrix $r\in\mathbb{R}^{3}$, and a rotation expressed by quaternion $q_{b}\in\mathbb{R}^{4}$. We limit the max radii of each part with a hyperparameter. To initialize these neural bones, we use the K-means algorithm provided by Scikit-learn (Pedregosa et al., 2011) to cluster 3D Gaussians representing the static self-model. We use cluster centers to initialize center points $p_{b}$. The rotation of each bone is initialized to $(0,0,0,1)$, and the radii are initialized to half of their max value. The number of bones can be either a custom value or one plus the number of joints of the robot. All parameters of neural ellipsoid bones are automatically optimized when training. We use a kinematic network to map joint angles into bone-specific transformation, which will be explained below. In the experiment, we found that assuming all neural bones have similar volumes can get the best result. In practice, we use eight bones for the 7-DOF robot in a simulation environment and seven for the 4-DOF robot in the real world. The 4-DOF robot we use has a large base, so we use multiple bones to represent it.
3.3 Kinematic network
The kinematic network predicts how bones transform from the rest pose (canonical space) to a joint configuration pose (observation space). The structure of our kinematic network is shown in Figure. 2. Assuming there are $M$ bones, the input joints are first encoded to bone-specific code $z_{m}$ with an encoder and then decoded to rotation and translation with MLP decoder $\Psi$. We follow the idea of Tertikas et al. (2023) to build our encoder. The joint angle vector $\theta$ is first encoded with an MLP encoder $f$ and then decomposed to per-bone codes with a multi-head attention transformer $\tau$. The whole encoding process is as follows:
$z_{m}=\tau(f(\theta)),m\in{1,2,\ldots,M}.$ (4)
Then the rotation $\Delta R_{m}^{\theta}$ and translation $\Delta t_{m}^{\theta}$ of bone $m$ at joint configuration $\theta$ are decoded from the bone-specific code by the decoder $\Psi$. Using $R_{m}^{c}$ to represent the rotation matrix derived from the quaternion of bone $m$ at rest pose and using $p_{m}^{c}$ to represent the center point of bone $m$ at rest pose, the rotation matrix $R_{m}^{\theta}$ and position $p_{m}^{\theta}$ at joint configuration $\theta$ are given by:
$R_{m}^{\theta}=\Delta R_{m}^{\theta}R_{m}^{c},$ (5)
$p_{m}^{\theta}=\Delta t_{m}^{\theta}+p_{m}^{c}.$ (6)
3.4 3D Gaussian deformation
We use linear blender skinning (LBS) (Sumner et al., 2007) to calculate the transformation of each 3D Gaussian based on the bone transformation that we get from the kinematic network. We use $w_{im}^{\rightarrow}$ to represent the skinning weight of 3D Gaussian $i$ with respect to bone $m$ when transforming from rest pose to joint configuration pose. Then, the center point $\mu_{i}^{\theta}$ and the rotation matrix $R_{i}^{\theta}$ derived from the quaternion of the deformed 3D Gaussian $i$ at joint configuration $\theta$ can be computed by:
$\mu_{i}^{\theta}=\sum_{m\in M}{w_{im}^{\rightarrow}(\Delta R_{m}^{\theta}(\mu_{i}^{c}-p_{m}^{c})+p_{m}^{c}+\Delta t_{m}^{\theta})}$ (7)
$R_{i}^{\theta}=\sum_{m\in M}{w_{im}^{\rightarrow}{\Delta R}_{m}^{\theta}}R_{i}^{c}$ (8)
where $\mu_{i}^{c}$ is the center point of 3D Gaussian $i$ at the rest pose, and $R_{i}^{c}$ is the rotation matrix of 3D Gaussian $i$ at rest pose. To simplify the expression of the transformation, we use $\mu_{i}^{\theta}=\Theta^{\theta,\rightarrow}(\mu)$ to represent (7). Similarly, we also define an inverse transformation $\mu=\Theta^{\theta,\leftarrow}(\mu_{i}^{\theta})$ that is calculated as follows:
$\mu_{i}^{c}=\sum_{m\in M}{w_{im}^{\leftarrow}((\Delta R_{m}^{\theta})^{-1}(\mu_{i}^{\theta}-p_{m}^{\theta})+p_{m}^{\theta}-\Delta t_{m}^{\theta})}.$ (9)
The calculation of skinning weights $w_{im}$ includes two parts. The original skinning weight $w_{im}^{o}$ is determined by the Mahalanobis distance between 3D Gaussian $i$ and bone $m$ , as mentioned by Yang et al. (2021):
$w_{im}^{o}=(\mu_{i}-p_{m})^{T}(R_{m})^{T}r_{i}R_{m}(\mu_{i}-p_{m})$ (10)
where $r_{i}$ is a $3\times 3$ diagonal radius matrix of the 3D Gaussian $i$. When computing the weights for the transformation from the rest pose to the joint configuration pose, the distance is computed with the center points and rotation matrices of 3D Gaussian $i$ and bone $m$ at rest pose. The weights for the inverse transformation are computed with the same values at the joint configuration pose.
Similar to the approach proposed by Yang et al. (2022), delta skinning weights are used for fine geometry. Delta skinning weight $w_{im}^{\Delta}$ is given by a weight MLP $\omega$:
$w_{im}^{\Delta}=\omega(\mu_{i},\theta)$ (11)
where $\theta$ is the joint angle vector. To calculate the forward transformation weights, $\theta$ is the rest pose joint angle vector. To calculate the inverse transformation weights, $\theta$ denote the joint configuration vector.
The final skinning weight is the sum of the two parts of weights and normalized by a softmax function:
$w_{im}=softmax(-s(w_{im}^{o}+w_{im}^{\Delta}))$ (12)
where $s$ is a learnable scaling factor.
With the 3D Gaussian deformation method, we can control the static self-model. Combining the static self-model and the deformation method, we build a dynamic and controllable self-model of a robot and obtain its morphology, kinematics, and texture.
3.5 Optimization
We observed that training 3D Gaussians, the kinematic network, and neural bones simultaneously often led to convergence issues. Additionally, due to the robot’s high degrees of freedom, the kinematic network struggled to train effectively with random joint configurations. To address these challenges, after training the static 3DGS self-model, we freeze the parameters of all 3D Gaussians and focus solely on training the kinematic network. During the initial phase of training, we restrict the robot’s joint motions to a limited range by using the small-range dataset and gradually relax these constraints over time. Once the range of motion reaches its maximum, we resume joint training of the 3D Gaussian parameters and the kinematic network. In total, we train the self-model over 600,000 steps.
3.6 Loss functions
Our optimization objective $L$ is the sum of several loss functions. We use the same image reconstruction loss $L_{render}$ in 3DGS, which is a combination of $L_{1}$ and D-SSIM to compare the input with the rendering image.
$L_{render}=(1-\lambda)L_{1}+\lambda L_{D-SSIM}.$ (13)
We also use a mask loss $L_{mask}$ that compares the foreground mask of the input image and the rendering foreground mask. $L_{mask}$ is computed by $L_{1}$ loss. Following previous work (Prokudin et al., 2023), we use the as-isometric-as-possible constrain (Kilian et al., 2007) to restrict the distance between neighboring 3D Gaussians to keep similar before and after deformation:
$L_{isopos}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}\left|d(\mu_{i}^{c}-\mu_{k}^{c})-d(\mu_{i}^{\theta}-\mu_{k}^{\theta})\right|\ }$ (14)
where $N$ is the total number of 3D Gaussians, $\mathcal{N}$ is the k-nearest neighborhood of 3D Gaussian $i$ We also use the local-rigid loss $L_{rigid}$ and the rotational loss $L_{rot}$ that is mentioned by Luiten et al. (2024) to constrain the kinematic network. Each of these losses is assigned a weight determined by the distance between 3D Gaussians:
$w_{i,k}=exp(-\lambda_{w}\|\mu_{k}^{c}-\mu_{i}^{c}\|_{2}^{2}).$ (15)
Then, the local-rigid loss $L_{rigid}$ is defined as follows:
$L_{rigid}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}w_{i,k}\|(\mu_{k}^{\theta}-\mu_{i}^{\theta})-R_{i}^{\theta}{(R_{i}^{c})}^{-1}(\mu_{k}^{c}-\mu_{i}^{c})\|_{2}}$ (16)
where $K$ is the number of k-nearest neighbors. The rotational loss is similarly given by:
$L_{rot}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}w_{i,k}\|R_{k}^{\theta}{(R_{k}^{c})}^{-1}-R_{i}^{\theta}{(R_{i}^{c})}^{-1}\|_{2}}.$ (17)
We use a 3D cycle loss as mentioned by Yang et al. (2022) to further constrain the kinematic network:
$L_{cycle}=\sum_{i=1}^{N}{\|\Theta^{\theta,\leftarrow}(\Theta^{\theta,\rightarrow}(\mu_{i}^{c}))-\mu_{i}^{c}\|_{2}^{2}}.$ (18)
We use a bone mask loss $L_{bone}$, a bone center loss $L_{center}$ , and a bone volume loss $L_{volume}$ to train neural bones. We treat each neural bone as a 3D Gaussian and render them into a binary image. The bone mask loss is given by the $L_{1}$ loss of the bone binary image and the robot’s mask. To discourage neural bones from being concentrated in a small region of the robot, we penalize small distances between their centers, as proposed by Noguchi et al. (2022):
$L_{center}=\frac{1}{M^{2}}\sum_{m\neq n}{\exp(\frac{\left|p_{m}-p_{n}\right|_{2}^{2}}{2\sigma^{2}})}$ (19)
where $\sigma$ controls the scale of the distances to be regularized. We also encourage that bones have comparable volume by the bone volume loss:
$L_{volume}=\frac{1}{M(M-1)}\sum_{i=1}^{M}{\sum_{k=1}^{i}{\left|\mathcal{V}(r_{i})-\mathcal{V}(r_{k})\right|}}$ (20)
where $\mathcal{V}(\cdot)$ means the volume of the neural ellipsoid bone calculated by the radius scaling matrices.
The final loss is a weighted sum of all losses mentioned above:
$L=$ $\lambda_{1}L_{render}+\lambda_{2}L_{mask}+\lambda_{3}L_{isopos}$ (21)
$+\lambda_{4}L_{rigid}+\lambda_{5}L_{rot}+\lambda_{6}L_{cycle}$
$+\lambda_{7}L_{bone}+\lambda_{8}L_{center}+\lambda_{9}L_{volume}$
where $\lambda_{i}$ (with $i=1,2,\cdots,9$) is the weight coefficient.
4 Results
We validated the proposed self-modeling method in a simulation environment using a 7-degrees-of-freedom (7-DOF) Franka robot and in an experimental setting with a 4-DOF OpenManipulator robot. Subsequently, we demonstrated the potential of our method in downstream tasks through several examples of motion planning and inverse kinematics.
Figure 3: Self-modeling result and ground truth in the simulation environment. Given a random joint configuration, we rendered the self-model to images from 3 different view angles. The joint configurations are shown in radians.
4.1 Simulation
To train the model, we followed the steps below to generate $400\times 400$ resolution images of the 7-DOF Franka robot in the Pybullet simulator (Coumans and Bai, 2016–2021). First, all joints of the robot were set to 0 radians, i.e., $\theta=[0,0,0,0,0,0,0]^{T}$ rad. A total of 100 images of the stationary robot were captured from various random view angles. Next, we actuated the robot by selecting different joint angles. Initially, we constrained all joint angles within $[-\frac{\pi}{6},\frac{\pi}{6}]$ and then expanded them to $[-\frac{\pi}{3},\frac{\pi}{3}]$. Finally, we limited the joint movement to the range of $[-\frac{\pi}{2},\frac{\pi}{2}]$. In most of our simulation, we maintained the joint angles in the final range, as this range allows the robot to perform a large number of tasks while simplifying the self-model training process. We took one image for each pose of the robot from a random view angle, and the camera parameters for each image were calculated and saved during the image generation process. A total of 10,000 images were generated: 500 images with joints constrained between $[-\frac{\pi}{6},\frac{\pi}{6}]$, another 500 images with joints constrainted between $[-\frac{\pi}{3},\frac{\pi}{3}]$, and the remaining 9,000 images with joints constrainted between $[-\frac{\pi}{2},\frac{\pi}{2}]$. Additionally, we captured 9,000 images without any joint constraints to demonstrate that our method can handle this situation effectively.
The results in the simulation environment are shown in Figure. 3. Four different joint configurations were randomly generated to control the self-model, and the self-model is rendered to images from three different views, including two horizontal views and a top view. Comparing the rendered images with ground truth images, it is obvious that the self-model built with our method correctly reflects the robot’s morphology, kinematics and texture under different joint configurations.
As illustrated in Figure. 4a, we assigned a distinct color to each neural bone and rendered them into an image, demonstrating the self-model’s ability to learn the robot’s kinematics. Eight neural bones were used in the self-model. They are not constrained to be linked together. The order of connection and the size of each bone are automatically determined by the self-model.
Figure 4: a Bone images in different joint configurations and the corresponding rendered images. Every neural bone is assigned a color randomly and rendered to images in the same way as in 3DGS. These neural bones form the kinematic link of the robot. The bone images are identical to the robot images at each joint configuration. b Simulation results when joint ranges are not limited. The reconstruction quality is worse than the case with limited joint ranges. However, the self-model still gives reliable results on how the robot moves at different joint configurations.
The quantitative results of the above four different joint configurations are presented in Table 1. The metrics we used are peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018), which are commonly used to evaluate synthetic image quality. The quantitative results for each joint configuration were calculated using seven images, consisting of six horizontal views around the robot and one top view. We also validated the proposed method in a test set containing 2,000 images captured from random view angles around the robot, with varying joint configurations. The results of the test set indicate that our self-model can correctly synthesize images of the robot at different joint configurations and view angles. It is worth noting that larger changes in joint configurations make it more challenging to obtain an accurate self-model and the test set includes joint configurations with minor changes. Consequently, the results for the 4 example joint configurations are worse than those of the test set.
Table 1: Quantitative results in simulation environment
Metric Configuration $\theta_{1}$ Configuration $\theta_{2}$ Configuration $\theta_{3}$ Configuration $\theta_{4}$ Test Set
[…]
We also present results without joint limits when posing the robot, as shown in Figure 4b. For each joint configuration, we display two images from different random view angles, with at least one of the robot arm’s seven joints exceeding $\frac{\pi}{2}$. Self-modeling with a larger range of joint angles is significantly more challenging, leading to results that are worse than those obtained within a smaller range. The robot’s end effector is not precisely reconstructed due to its position at the end of the kinematic chain and the complexity of its movement. However, other parts of the robot are still accurately modeled. Compared to the method proposed by Schulze and Lipson (2024), which uses a NeRF-based method for self-modeling, our method reconstructs the robot’s shape more precisely, allowing easier identification of the robot’s texture and morphology in the rendered images.
We also present results without joint limits when posing the robot, as shown in Figure 4b. For each joint configuration, we display two images from different random view angles, with at least one of the robot arm’s seven joints exceeding $\frac{\pi}{2}$. Self-modeling with a larger range of joint angles is significantly more challenging, leading to results that are worse than those obtained within a smaller range. The robot’s end effector is not precisely reconstructed due to its position at the end of the kinematic chain and the complexity of its movement. However, other parts of the robot are still accurately modeled. Compared to the results presented by Schulze and Lipson (2024), which uses a NeRF-based method for self-modeling, our method reconstructs the robot’s shape more precisely, allowing easier identification of the robot’s texture and morphology in the rendered images.
4.2 Physical experiment
We utilized a smartphone and an Omnivision OV5640 color sensor attached to the end effector of a robot to capture images of the OpenManipulator robot. All images were resized to a resolution of $640\times 480$ while maintaining the aspect ratio. Following a procedure similar to that in the simulation, we first used both cameras to capture images with the robot in its zero pose. The smartphone was free to capture images from various view angles, whereas the color sensor could only capture images from fixed angles. To simplify data collection, we captured images in a forward-facing manner instead of a $360^{\circ}$ inward approach. Notably, multiple fixed cameras could be used in place of the robot’s color sensor during data collection. Camera parameters were computed using structure-from-motion (SfM) (Schönberger and Frahm, 2016) based on the static scene. Subsequently, we randomly posed the robot and used only the color sensor to capture images from the calculated view angles. We restricted the joint angles between $[-\frac{\pi}{6},\frac{\pi}{6}]$ to capture 100 images, then posed the robot without joint limitations to capture an additional 1,000 images. We employed the segment anything model (SAM) (Kirillov et al., 2023) to generate masks and segment the robot from the background.
Figure 5: Physical experiments on the OpenManipulator robot with no joint limitation. All joint configurations were randomly generated, and the images were rendered from novel view angles. Joint configurations are shown in radians.
The results of physical experiments are shown in Figure 5. We randomly selected three sets of joint angles that were not part of the training data, rendered the corresponding robot images using the trained model, and compared them to the ground truth. The comparison reveals that the proposed method accurately represents the robot’s morphology and texture, further validating the effectiveness of our approach.
Conducting self-modeling with real-world data is more challenging than with simulation data. Thus, the real-world results are not as accurate as those obtained from simulations. This can be attributed to several factors: (1) Potential inaccuracies in the camera parameters calculated by SfM, (2) Imperfect segmentation of the robots from the background in certain images, and (3) A smaller volume of data is available in real-world experiments compared to simulation. Nevertheless, our real-world results are significantly closer to the ground truth compared to previous work by Chen et al. (2022).
4.3 Applications
In this section, we use the proposed self-model to complete some downstream tasks: reaching a target and reaching a target while avoiding an obstacle. We also show that our self-model can solve the inverse kinematics problem.
4.3.1 Reaching a target
A green ball was placed within the robot’s workspace, requiring the robot to use its self-model to determine how to make contact with the ball using a specific part of its body. Our neural bone representation simplifies identifying specific parts of the robot and enables targeted control to reach the desired object. We use the robot’s end effector as an example. Starting from a resting configuration, the joint angles are progressively adjusted to minimize the distance between the target and the center of the neural bone associated with the end effector. This is accomplished by fixing the self-model parameters and applying gradient descent to automatically optimize the joint angles. In practice, the AdamW (Loshchilov, 2017) optimizer is utilized, and the loss function is defined as the mean squared error (MSE) between the bone center and the target position. Unlike method of Chen et al. (2022), by altering the neural bone applied in the loss function, the robot can be controlled to reach the target using different parts without the need for additional training.
4.3.2 Reaching a target while avoiding an obstacle
An obstacle was placed at a location the robot would pass through after completing the first task of reaching the target. The robot is required to find a new joint configuration to reach the target, as the original configuration would result in a collision with the obstacle. To address this, an additional loss function is introduced alongside the MSE loss function used in the previous task. All 3D Gaussians are treated as point clouds, and the distances $D$ between the points and the obstacle are computed. The loss function is then formulated as:
$L=\sum_{i=1}^{N}{\max(0,\varepsilon-D_{i})}$ (22)
where $\varepsilon$ is a threshold that controls how close the robot can get to the obstacle. In our experiment, about 50,000 3D Gaussians were used in the self-model. All 3D Gaussians were forced to keep a distance from the obstacle.
The two applications above are demonstrated in Figure 6. Images in the first row show how our self-model controls the end effector of the robot to reach the target, shown here as a green ball. The robot’s final pose in this task is shown in the fifth image of the first row (Movie S1). Next, an obstacle was placed at the position of the robot’s fourth link in this final pose, and images in the second and third rows show two perspectives of our self-model guiding the robot to complete the target-reaching task while avoiding the obstacle. Clearly, the robot adopts a new pose to avoid the obstacle and reach the target (Movie S2 and Movie S3). The minimum distance between the obstacle and the robot’s 3D Gaussian representations is also shown in the last row of Figure 6b.
Figure 6: Results of applications. a Snapshots and joint angles of the robot with respect to step number. b Snapshots from different perspectives, joint angles of the robot, and the minimum distance between the obstacle and the robot with respect to step number. The unit of the joint angle is radian. At the beginning, all joints of the robot are set to 0. Using solely the self-model, the robot is able to find joint configurations that can fulfill the tasks. c Results of inverse kinematics estimation. The ground truth joints are [-0.204, -0.9319, -0.3448, -1.5012, 0.4474, 0.105, 0.3166]. The estimated joints are [-0.2127, -0.9321, -0.3051, -1.5050, 0.4863, 0.0258, 0.3113]. Joint angles in radians are slightly different, but the morphology of the robot is close to the ground truth.
4.3.3 Robot inverse kinematics estimation
With our self-model, we can estimate the robot’s inverse kinematics. Given images of the robot, our self-model can predict a joint configuration to pose the robot similarly. After freezing self-model parameters and treating joint angles as optimizable, we employ gradient descent to approximate the joint configuration. Comparing rendered and ground truth images, we optimize joint angles to achieve reasonable values. In practice, we use two images to estimate the robot’s pose. Our estimation results are illustrated in Figure 6c. Although estimated joint angles differ slightly from the actual angles, rendered images at the estimated joint angles match ground truth images. Our self-model effectively poses the robot in alignment with the ground truth. Additional images from varied view angles could further improve joint angle estimation accuracy.
5 Discussion
We propose a 3DGS-based robot self-modeling method that eliminates the need for depth information. This approach achieves high-quality modeling of the robot’s morphology, texture, and kinematics. Our method initiates modeling from the robot’s rest pose. We employ learnable neural ellipsoid bones of similar volume to represent morphology and texture. A kinematic network is trained to capture the robot’s kinematic properties. Each 3D Gaussian’s movement is governed by transformations of neural bones, enabling full model control. Rendering 3D Gaussians into images enables us to visualize the robot’s morphology and texture. Our self-model accurately reconstructs the robot across various joint configurations, achieving high-quality image synthesis.
The self-model developed using our method can be applied in downstream tasks. With this self-model, the robot is capable of solving motion-planning tasks. Our neural bone representation directly models the robot’s structure, allowing us to control different parts of the robot by specifying which bone to optimize in the loss function without requiring additional information or training to adjust the self-model. Since our self-model can synthesize high-quality images, it can also estimate the robot’s inverse kinematics by replicating input image poses.
However, our method can be further improved in several ways. (1) A better initialization of our neural bones is possible. The current method requires a predefined number of bones to initialize the structure. We also assume that each bone is similar in size, but in reality, the size of robot links may vary, necessitating multiple bones to represent a single link. (2) Our method may not always accurately reconstruct the robot at certain joint configurations, especially when the joint’s movement range is unrestricted. It is worth researching how to obtain a more accurate self-model across all joint configurations. (3) It is important to find a method to map the 3DGS self-model to a mesh model and modify it for compatibility with common robot control pipelines, such as MOVEIT (Coleman et al., 2014) or current reinforcement learning toolkits. (4) We need multiple views to reconstruct the model, and the number of views, along with the camera parameters, can affect the final result. Researching how to self-model the robot with high quality using fewer views and without known camera parameters is a valuable direction to explore.
6 Author Contributions
7 Statements and declarations
7.1 Ethical considerations
Not applicable.
7.2 Consent to participate
Not applicable.
7.3 Consent for publication
Not applicable. {dci} The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
{funding}
The author(s) disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported in part by the National Natural Science Foundation of China [grant number 62173352], and in part by the Guangdong Basic and Applied Basic Research Foundation [grant number 2024B1515020104].
{sm}
Movie S1: The robot is controlled by its self-model to reach the green target.
Movie S2: The robot is controlled by its self-model to reach the green target while avoiding the red obstacle at viewpoint 1.
Movie S3: The robot is controlled by its self-model to reach the green target while avoiding the red obstacle at viewpoint 2.